# The Green Peace Dividend: the Effects of Militarization on Emissions and the Green Transition

### v3

@@ Abstract
This paper argues that military buildups lead to a significant rise in greenhouse gas emissions and can disrupt the green transition. Identifying military spending shocks, I use local projections to show that a percentage point rise in the military spending share leads to a 0.9-2% rise in total emissions, as well as a 1% rise in emission intensity, depending on the economy’s overall emission intensity. Using a dynamic production network model calibrated for the US, I find that a permanent shock of the same size would increase total emissions by between 0.36% and 1.81%, and emission intensity by between 0.22% and 1.5%. The empirical analysis indicates that green patenting is reduced by 10-25% following such a shock, and the model suggests that investment in renewables could be crowded out by defence spending, hindering the energy transition. These effects can significantly raise climate damages, temperatures, and can lead to a quicker exhaustion of the carbon budget.
@@ 1 Introduction
This paper aims to uncover, measure, and understand the effects of militarization on climate change mitigation. In particular, I investigate two main channels. First, military buildups can raise greenhouse gas emissions directly, through increased energy and fuel use and a larger volume of military procurements. Armed forces need more energy and fuel to run their military bases, planes, bombers, tanks, etc. at a higher intensity, which increases energy use, but they also need more of these weapons. Defence manufacturing is an energy-intensive part of the economy, but especially its inputs, such as primary metals, rubber, or chemicals are very energy-intensive. Hence, increasing demand for these goods can drive up the emission intensity of an economy and consequently total emissions.
Second, military buildups can impede the green transition through indirect economic effects. Such a demand shock can lead to a crowd-out in green investment and innovation. Defence and fossil fuel industries will require large amounts of investment to expand capacity to the level where it can supply a substantially more active military. This can lead to a scarcity of investment funds, delaying and potentially disrupting the financing of the green transition.
@@
The result that the risk of military conflict can worsen climate change is important on its own because it shows that there is a "green peace dividend" - an additional cost of conflict. In section 5, I show that the measured effects can significantly raise climate damages, increase temperatures, and lead to the quicker exhaustion of the carbon budget. However, we also know that climate change can increase the likelihood of conflicts (see, e.g., Burke, Hsiang, and Miguel (2015) for a good summary of this literature). Hence, a positive feedback loop between conflict and climate change can increase the destructiveness of both.
@@
Ramey (2011), Ramey and Shapiro (1998), Auerbach and Gorodnichenko (2012), Hall (1980), Barro (1981), Nakamura and Steinsson (2014), and Miyamoto et al. (2019) exploit military spending shocks to measure the effect of government spending on different economic outcomes - however, they do not analyse the link between military spending and emissions. Couttenier et al. (2022) and Korovkin and Makarin (2022) use production network models to estimate the economic effects of wars, in the context of the Maoist insurgency in India and the Ukraine war, respectively. I use methods similar to these papers to understand the effect of military activity, but I focus on climate-relevant impacts.
@@ 2.1.1 Data
In my empirical analysis, I use annual GHG emissions data from EDGARv8.0. I also use data from the Correlates of War (National Material Capabilities version 6.0) for energy consumption in tons of coal equivalent, as well as for military expenditures, military personnel, crude steel production, population, and urban population. National accounts data is from the Penn World Tables (version 10.01). I’m using data from V-Dem v13 measuring how close a given country is to the "ideal of electoral democracy" (Coppedge et al, 2023). I use OECD data for total and environmental patents.
The data covers the period 1970-2016. In my main sample, I include 20 NATO countries: Belgium, Bulgaria, Canada, Denmark, Finland, France, Germany, Greece, Hungary, Italy, the Netherlands, Norway, Poland, Portugal, Romania, Spain, Sweden, Turkey, the United Kingdom, and the United States. These are countries that are likely to have good data for the variables of interest in the period analysed. They also yield a balanced panel in my analysis (see below). In addition, the effect of military buildups in these countries - NATO members - is of special interest following the Russian invasion of Ukraine. Many of them are also countries that have some arms industry, which raises the domestic absorption of the military budget (I discuss below why this is important).
@@ 2.1.2 Method
To extract the causal effect of military spending on the outcomes of interest, I follow the approach in Bilal and Känzig (2024) and first extract military spending shocks using a Hamilton (2018) filter, after which I estimate the dynamic effect of such a shock on the outcomes of interest using panel local projections (Jordà (2005), Jordà et al. (2020)). I assume that these shocks are exogenous, as sudden changes in defence spending are most likely to follow strategic, and not economic, considerations. It is common in the literature to assume that military spending shocks occur due to geopolitical events, and not due to endogenous economic considerations. For example, Miyamoto et al. (2019) assume that "changes in military spending are exogenous to economic conditions". A similar argument can be found in Hall (1980).
@@ 3 Model
This section describes the theoretical model. I use it to provide additional evidence on the effect of a military buildup on greenhouse gas emissions, to understand what other channels it impacts the green transition through, as well as the mechanisms through which these effects work.
@@ 3.1 Setup
Where $E$ denotes total emissions, $\epsilon_{i}$ denotes the share of industry i’s scope 1 emissions191919”Scope 1 emissions are direct greenhouse (GHG) emissions that occur from sources that are controlled or owned by an organization (e.g., emissions associated with fuel combustion in boilers, furnaces, vehicles)” (US EPA) out of the total emissions of the economy, $\epsilon_{HH}$ denotes the share of household scope 1 emissions out of total emissions, and $\epsilon_{Govt}$ is the share of the government’s scope 1 emissions out of the total.
@@ 3.3.1 Permanent shock
Table 4: Change in industry-level output in response to bringing the military expenditure share to its Korean War peak level.
@@ 4 The Green Peace Dividend and policy implications
I use three main methods to assess the impacts of the measured effects on GHG: by using the social cost of carbon (SCC), by using the carbon budget, and by looking at the global path of surface temperatures.
@@
First, focusing solely on these direct effects of military buildups on GHG emissions, it is possible to translate the damages due to induced emissions into dollar terms using the social cost of carbon (SCC). For example, if one uses the social cost of carbon (SCC) provided by the EPA (2023), equal to \$190 per ton of CO2 at a near-term discount rate of 2%, it is possible to consider a scenario in which the US permanently doubled its military spending share in 2017 and to calculate the climate damages resulting from this policy change using the SCC. Total GHG emissions in the US in 2017 were equal to 6.09 Gt of CO2e. The model estimate of a 1.18-6% rise in emissions in response to such a shock (with a military spending share of 3.3% in 2017) leads to global climate damages equivalent to approximately 13.6-69.4 billion dollars per year - or approximately 0.07-0.35% of 2017 US GDP. [^32] This calculation ignores the transition path to the new steady-state, as well as the fact that emissions change from their 2017 level for reasons independent of defence policy. More recent - and larger - estimates, such as the SCC calculated by Bilal and Känzig (2024), equal to \$1056 per ton of CO2, would yield an interval of 75.9-385.9 billion dollars per year - or approximately 0.39-2% of 2017 US GDP. Note that these simple calculations ignore the convexity of climate damages.
Second, one can reframe the consequences in the following way: rising GHG emissions due to military buildups lead to the quicker exhaustion of our carbon budget. [^33] The amount of GHG emissions we can afford before breaching a given threshold of warming, such as 2 degrees Celsius. Hence, militarization would reduce our chances of avoiding any given global warming limit. In particular, according to the Global Carbon Budget 2023 (Friedlingstein et al. (2023)), the remaining carbon budget for a 50% likelihood to limit global warming to 2 °C is equal to 1150 Gt of CO2 from the beginning to 2024. Under 2023 emission levels (approximately 40.9 Gt of CO2), this would be exhausted in 28 years. Now, assuming that all countries permanently raise their military expenditure share by 3.3 percentage points (as we assumed in the SCC calculations for the US), and that the effects are similar to the US (emissions rise by 1.18-6%), then the carbon budget could be exhausted 0.22-1.5 years earlier.
@@
Third, one can check how a global military buildup of 3.3 percentage points would change the path of temperatures, under different shared socioeconomic pathways (SSP) (Riahi et al. (2017)). Here, I use the central estimate of the effect of a military expenditure shock on emissions, which implies that a permanent shock of 3.3 p.p. raises GHG emissions by 3%. Using live.magicc.org, an online software relying on MAGICC7 (Meinsshausen et al. (2020)) (which is an integrated assessment model), I calculate that the difference would be minimal under the SSP1-1.9 scenario (which assumes that GHG emissions are very low, and are reduced to zero by 2050), with temperatures being higher by approximately 0.01 °C in 2100 (see figure 17). [^34] In the simulations, I set CO2 emissions from fossil fuel use and industrial activity to be higher by 3%, after 2015. However, under the SSP3 scenario (where GHG emissions double by 2100), temperatures would be higher by 0.05 °C by 2100, indicating that the effect of a military buildup becomes worse under worse scenarios. Under the SSP5 scenario, temperatures would be higher by more than 0.06 °C. [^35] For an explanation of the different shared socioeconomic pathways, see: https://www.carbonbrief.org/explainer-how-shared-socioeconomic-pathways-explore-future-climate-change/ This indicates that under progressively worse scenarios, the effect of a military buildup on temperatures also becomes stronger, as a 3% rise in a higher baseline amount of emissions leads to a larger change in the level of emissions, leading, in turn, to a larger change in the level of the temperature. Note that temperatures in 2100 rise with the shown scenarios (so they are highest under SSP5, and lowest under SSP1 - 1.9), and temperatures peak before 2100 only under the SSP1 scenarios.

### v4

@@ Abstract
This paper argues that military buildups lead to a significant rise in greenhouse gas emissions and can disrupt the green transition. Identifying military spending shocks, I use local projections to show that a percentage point rise in the military spending share leads to a 0.9-2% rise in total emissions, as well as a 1% rise in emission intensity, depending on the economy’s overall emission intensity. Using a dynamic production network model calibrated for the US, I find that a permanent shock of the same size would increase total emissions by between 0.36% and 1.81%, and emission intensity by between 0.22% and 1.5%. The empirical analysis indicates that green patenting is reduced by 10-25% following such a shock, and the model suggests that investment in renewables could be crowded out by defence spending under certain circumstances, hindering the energy transition. These effects can significantly raise climate damages and temperatures. Depending on the social cost of carbon and the composition of a military spending shock, I estimate that doubling the military spending share in the US in 2017 (equal to 3.3% of GDP) would have led to climate damages equivalent to between 0.07-2.6% of GDP per year.
@@ 1 Introduction
This paper aims to uncover, measure, and understand the effects of militarization on climate change mitigation. In particular, I investigate two main channels. First, military buildups can raise greenhouse gas emissions directly, through increased energy and fuel use and a larger volume of military procurements. Armed forces need more energy and fuel to run their military bases, planes, bombers, tanks, etc. at a higher intensity, which increases energy use. They also need more of these weapons, and defence manufacturing is an energy-intensive part of the economy, but especially its inputs, such as primary metals, rubber, or chemicals are very energy-intensive. Hence, increasing demand for these goods can drive up the emission intensity of an economy and consequently total emissions.
Second, military buildups can impede the green transition through indirect economic effects. Such a demand shock can lead to a crowd-out in green investment and innovation. Defence and fossil fuel industries will require large amounts of investment to expand capacity to the level where it can supply a substantially more active military. This can lead to a scarcity of investment and innovation funds for climate change mitigation and adoption, delaying and potentially disrupting the green transition.
@@
The result that the risk of military conflict can worsen climate change is important on its own because it shows that there is a "green peace dividend" - an additional cost of conflict. In section 5, I show that the measured effects can significantly raise climate damages and increase temperatures. The results from the model indicate that if the military spending share was doubled in the US in 2017 (when it was 3.3% of GDP), damages due to higher emissions would equal between 0.07-2.6% of GDP per year, depending on the social cost of carbon and the composition of the military spending shock. This calculation does not take into account potential effects on green innovation and investment.
However, we also know that climate change can increase the likelihood of conflicts (see, e.g., Burke, Hsiang, and Miguel (2015) for a good summary of this literature). Hence, a positive feedback loop between conflict and climate change can increase the destructiveness of both. I leave the evaluation of this feedback loop to future work.
@@
Ramey and Shapiro (1998), Ramey (2011), Ilzetzki et al. (2013), Auerbach and Gorodnichenko (2012), Hall (1980), Hall (2009), Barro (1981), Barro and Redlick (2011), Gordon and Krenn (2010), Nakamura and Steinsson (2014), and Miyamoto et al. (2019) exploit military spending shocks to measure the effect of government spending on different economic outcomes - however, they do not analyse the link between military spending and emissions. Couttenier et al. (2022) and Korovkin and Makarin (2022) use production network models to estimate the economic effects of wars, in the context of the Maoist insurgency in India and the Ukraine war, respectively. I use methods similar to these papers to understand the effect of military activity, but I focus on climate-relevant impacts.
@@ 2.1.1 Data
I use annual GHG emissions data from EDGARv8.0. I also use data from the Correlates of War (National Material Capabilities version 6.0) for energy consumption in tons of coal equivalent, as well as for military expenditures, military personnel, crude steel production, population, and urban population. National accounts data is from the Penn World Tables (version 10.01). I’m using data from V-Dem v13 measuring how close a given country is to the "ideal of electoral democracy" (Coppedge et al, 2023). I use OECD data for total and environmental patents.
The data covers the period 1970-2016, except for the data on patents, which covers 1978-2016. In my main sample, I include 20 NATO countries: Belgium, Bulgaria, Canada, Denmark, Finland, France, Germany, Greece, Hungary, Italy, the Netherlands, Norway, Poland, Portugal, Romania, Spain, Sweden, Turkey, the United Kingdom, and the United States. These are countries that are likely to have good data for the variables of interest in the period analysed. They also yield a balanced panel in my analysis (see below). In addition, the effect of military buildups in these countries - NATO members - is of special interest following the Russian invasion of Ukraine. Many of them are also countries that have some arms industry, which raises the domestic absorption of the military budget (I discuss below why this is important).
@@ 2.1.2 Method
To extract the causal effect of military spending on the outcomes of interest, I follow the approach in Bilal and Känzig (2024) and first extract military spending shocks using a Hamilton (2018) filter, after which I estimate the dynamic effect of such a shock on the outcomes of interest using panel local projections (Jordà (2005), Jordà et al. (2020)). I assume that these shocks are exogenous, as sudden changes in defence spending are most likely to follow strategic, and not economic, considerations. It is common in the literature to assume that military spending shocks occur due to geopolitical events or strategic considerations, and not due to endogenous economic considerations. For example, Miyamoto et al. (2019) assume that "changes in military spending are exogenous to economic conditions". A similar argument can be found in Hall (1980).
@@ 2.2 Results
2.3 Robustness checks
Alternative transformations of military spending
One concern which could arise with regard to the shock identification procedure is that the military expenditure share, $M_{t}=m_{t}/Y_{t}^{N}$ (where $m_{t}$ is nominal military expenditure at time t, and $Y_{t}^{N}$ is nominal GDP) depends on a potentially endogenous factor - nominal output - even if military expenditures are exogenous, as argued before. Concretely, it would be problematic if I were capturing recessions and not military buildups (that is, drops in the denominator instead of rises in the numerator).
In order to remove this potential source of endogeneity, I consider two alternative formulations. First, instead of using the military expenditure share, I use a Gordon-Krenn transformed military expenditure variable (Gordon and Krenn (2010)). That is, I use:
$M^{GK}_{t}=\frac{m_{t}^{*}}{Y^{*}_{t}}$
Where $m_{t}^{*}$ is nominal military spending divided by the GDP deflator, while $Y^{*}_{t}$ is a measure of "potential GDP" - in my case, I set this equal to the quadratic trend of real GDP, following Ramey (2016). Unlike actual nominal GDP, potential GDP does not vary much and does not depend on the business cycle, and hence is unlikely to be endogenous in our setting.
I extract shocks to the Gordon-Krenn transformed military spending share, $M^{GK}_{t}$, using the Hamilton (2018) method as before, and run the local projections with the baseline NATO sample and specification. The results can be seen on figure 13 - instead of a 1 p.p. rise in the military spending share, the shock can now be interpreted as a 1 p.p. rise in the share of real military spending in potential GDP. Clearly, the results are extremely similar to the baseline results on figure 4.
Figure 13: Gordon-Krenn transformed military expenditures - emissions and energy
Note: the figure shows the IRFs of total emissions, energy use, emission intensity, and energy intensity in response to a military spending shock corresponding to a 1 p.p. rise in the Gordon-Krenn military spending share. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
One can also use the Hall-Barro-Redlick transformation (following Hall (2009) and Barro and Redlick (2011)) instead of the Gordon-Krenn transformation to get rid of endogeneity. The Hall-Barro-Redlick transformed variable gives the change in real military spending as a share of year $t-1$ real GDP:
$M^{HBR}_{t}=\frac{m_{t}^{*}-m_{t-1}^{*}}{Y^{R}_{t-1}}$
Where $Y^{R}$ denotes real GDP, and $m^{*}$ is as before. $M^{HBR}_{t}$ depends only on the last period’s real output, and is unrelated to the current period’s military spending shock. [^14] Note that I control for period $t-1$ changes in GDP per capita. Then, I use $M^{HBR}_{t}$ directly as a shock in my local projection formulation. The results can be seen on figure 14. Now, the shock should be interpreted as a 1 percent rise in real military spending relative to the last period’s real output. Again, the results are virtually unchanged compared to the baseline results on figure 4.
Figure 14: Hall-Barro-Redlick transformed military expenditures - emissions and energy
Note: the figure shows the IRFs of total emissions, energy use, emission intensity, and energy intensity in response to a military spending shock corresponding to a 1 percent rise in military spending relative to last period’s real output. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
Strategic spillover effects
If a given country raises its military spending share due to an exogenous shock shift in policy, other countries might react by changing their own military expenditures. For example, strategic allies might reduce their own expenditures because they free ride on their allies’ military capabilities - or, alternatively, they might increase their own expenditures because they need to help their allies (e.g. in case of war). Strategic adversaries might react by raising their own military expenditures. All this means that the global effects of a military spending shock might not be equal to their domestic effects: if one country’s rearmament drive leads to the rest of the world also rearming, the global rise in emissions might be an order of magnitude larger than the domestic effect.
In my sample, there is enough data from 1970 on to look at the effect of US military spending shocks on its strategic allies, which I define as those countries that were members of NATO in 1982.151515Eastern European countries became allies after 1989 after having been adversaries in the previous decades - hence, I do not consider them in this exercise. That is, I look at the effects of US military expenditure shocks on the military expenditure share of Belgium, Canada, Denmark, France, Germany, Greece, Italy, the Netherlands, Norway, Portugal, Spain, Turkey, and the United Kingdom, using the original local projection setup (2), without time fixed effects (which would filter out the common shock). As the US was one of the two most important military powers in this sample, this is a reasonable setup - I do not have data on Russian military expenditures separate from the USSR before 1991.
Figure 15: Spillover effect of US shocks
Note: the figure shows the IRFs of the military spending share in 1982 NATO countries in response to a US military spending shock corresponding to a 1 percentage point rise in the military spending share. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
The results can be seen on figure 15. In response to a 1 p.p. rise in the US military spending share, NATO allies raise their military spending share by 0.1-0.2 percentage points in the first 4 years after the shock, following which it returns to what it would have been absent the shock. This suggests that the US’s allies generally "showed solidarity" with the US following spending shocks, in the sample. This would suggest that the effects of a military spending share shock in the US would have important positive spillover effects in its NATO allies, meaning that my results might underestimate the global effects.
I do not present results for the strategic adversaries of the US. The reason for this is that there are only two countries in the sample which can be reasonably categorized as adversaries of the US throughout the sample: China and Russia, and I do not have data on Russia’s military expenditures before 1990. It would be important for future work to estimate empirically spillover effects on strategic adversaries.
@@ 3 Model
This section describes the theoretical model. I use it to provide additional evidence on the effect of a military buildup on greenhouse gas emissions, to understand through what other channels it impacts the green transition, as well as the mechanisms through which these effects work.
@@ 3.1 Setup
Where $E$ denotes total emissions, $\epsilon_{i}$ denotes the share of industry i’s scope 1 emissions212121”Scope 1 emissions are direct greenhouse (GHG) emissions that occur from sources that are controlled or owned by an organization (e.g., emissions associated with fuel combustion in boilers, furnaces, vehicles)” (US EPA) out of the total emissions of the economy, $\epsilon_{HH}$ denotes the share of household scope 1 emissions out of total emissions, and $\epsilon_{Govt}$ is the share of the government’s scope 1 emissions out of the total.
@@ 3.3.1 Permanent shock
Table 4: Change in industry-level output in response to doubling or bringing the military expenditure share to its Korean War peak level.
@@ 4 The Green Peace Dividend and policy implications
I use two main methods to assess the impacts of the measured effects on GHG: by using the social cost of carbon (SCC), and by looking at the global path of surface temperatures.
@@
First, focusing solely on these direct effects of military buildups on GHG emissions, it is possible to translate the damages due to induced emissions into dollar terms using the social cost of carbon (SCC). For example, if one uses the social cost of carbon (SCC) provided by the EPA (2023), equal to \$190 per ton of CO2 at a near-term discount rate of 2%, it is possible to consider a scenario in which the US permanently doubled its military spending share in 2017 and to calculate the climate damages resulting from this policy change using the SCC. Total GHG emissions in the US in 2017 were equal to 6.09 Gt of CO2e. The model estimate of a 1.18-6% rise in emissions in response to such a shock (with a military spending share of 3.3% in 2017) leads to global climate damages equivalent to approximately 13.6-69.4 billion dollars per year - or approximately 0.07-0.35% of 2017 US GDP. [^34] This calculation ignores the transition path to the new steady-state, as well as the fact that emissions change from their 2017 level for reasons independent of defence policy. More recent - and larger - estimates, such as the SCC calculated by Bilal and Känzig (2024), equal to \$1367 per ton of CO2, would yield an interval of 98.2-500 billion dollars per year - or approximately 0.5-2.6% of 2017 US GDP. Note that these simple calculations ignore the convexity of climate damages.
@@
Second, one can check how a global military buildup of 3.3 percentage points would change the path of temperatures, under different shared socioeconomic pathways (SSP) (Riahi et al. (2017)). Here, I use the central estimate of the effect of a military expenditure shock on emissions from the model, which implies that a permanent shock of 3.3 p.p. raises GHG emissions by 3%. Using live.magicc.org, an online software relying on MAGICC7 (Meinsshausen et al. (2020)) (which is an integrated assessment model), I calculate that the difference would be minimal under the SSP1-1.9 scenario (which assumes that GHG emissions are very low, and are reduced to zero by 2050), with temperatures being higher by approximately 0.01 °C in 2100 (see figure 19). [^35] In the simulations, I set CO2 emissions from fossil fuel use and industrial activity to be higher by 3%, after 2015. However, under the SSP3 scenario (where GHG emissions double by 2100), temperatures would be higher by 0.05 °C by 2100, indicating that the effect of a military buildup becomes worse under worse scenarios. Under the SSP5 scenario, temperatures would be higher by more than 0.06 °C. [^36] For an explanation of the different shared socioeconomic pathways, see: https://www.carbonbrief.org/explainer-how-shared-socioeconomic-pathways-explore-future-climate-change/ This indicates that under progressively worse scenarios, the effect of a military buildup on temperatures also becomes stronger, as a 3% rise in a higher baseline amount of emissions leads to a larger change in the level of emissions, leading, in turn, to a larger change in the level of the temperature. Note that temperatures in 2100 rise with the shown scenarios (so they are highest under SSP5, and lowest under SSP1 - 1.9), and temperatures peak before 2100 only under the SSP1 scenarios.
@@ 5 Conclusion
My analysis, based on the model results, indicates that doubling the military spending share in the US in 2017 would have led to climate damages equal to between 0.07-2.6% of GDP per year, depending on the SCC and the shock composition. Importantly, these damage calculations do not take into account potential effects on green innovation or investments.